OpenStudy (boothepest1):

Alan is 5 years older than Bob. Bob is twice as old as Cindy. All their ages add up to 20. What are their ages? Alan is____ Bob is____ Cindy is____ I'm in highschool we are doing Throwback Thursday in Math. I feel silly but I'm too exhausted to think. XD

OpenStudy (wolf1728):

A = B +5 B = 2*C A + B + C =20 Can you solve it from here?

OpenStudy (marcoreus11):

^